Solve each system using the specified method. Elimination
5x-4y=-1
10x+2y=3

Answer by checkley77(12569)

5x-4y=-1
10x+2y=3 multiply this equation by 2 & add
20x+4y=6
5x-4y=-1
---------------------
25x=5
x=5/25
x=1/5 ans.
5*1/5-4y=-1
1-4y=-1
-4y=-1-1
-4y=-2
y=-2/-4
y=-1/2 ans.
Proof:
10*1/5+2*1/2=3
2+1=3
3=3